Bubba is working on the loft in the barn his farm one day. The loft is 3.54 meters tall. Bubba decides to throw a hammer downward with an initial downward velocity of 3.38 m/s in order to hit a mouse that he sees crawling across the floor of the barn.

a) How long does it take for the hammer to reach the ground?

b) How fast is the hammer moving when it hits the ground?
Answer #1

Given,

H = 3.54 m ; v = 3.38 m/s

a)We know that

t = sqrt (2h/g)

t = sqrt (2 x 3.54/9.81) = 0.85 s

Hence, t = 0.85 s

b)We know that

v = u + at

v = 3.38 + 9.8 x 0.85 = 3.38 + 8.34 = 11.72 m/s

v = 11.72 m/s